# Michael Callen Miguel Fajardo-Steinhäuser Michael G. Findley

### v2

@@ 5 Interpreting magnitudes of needs results
The Afghanistan Cash and Voucher Working Group (CVWG) estimates a Basic Food Basket for a family of seven costs approximately 96 USD per month (Bete, 2022), roughly equivalent of two biweekly direct aid transfers of 4000 AFA and a monthly survey incentive of 350 AFA. This basket was composed of 89 kg wheat flour, 21 kg domestic rice, 7 liters vegetable oil, 9 kg pulses, and 1 kg salt at prevailing exchange rates in August 2022. For reference, the Minimum Expenditure Basket including food, healthcare, shelter, and other components totaled almost twice as much at 181.36 USD. In mid-December 2022, we contacted 25 merchants serving beneficiaries to better understand spending patterns and realized prices. We confirmed that the most popular purchases were wheat flour, cooking oil and sugar, and using merchant-specific estimates of prices and volumes, verified that the cost of a typical basket of goods matched the total value of the aid payments and survey incentive.
It is natural to ask whether the estimated impacts on basic needs are in line with what we should expect given the size of payments – and more specifically, why hunger appears to persist throughout the length of the intervention. As noted above, we observe large reductions in each of the four primary food security outcomes, and these measures kept improving throughout the two months of payments (Table 2, Panel A). While not eliminated, the number of days in which a person skipped meals in the past week declined after each payment (Figure A2, Panel A). To better understand the explanatory factors for this persistence, we completed a brief survey on January 2 & 3, 2023 with 52 randomly selected treatment group participants who reported any skipped meals in the second month of surveys (ie. December 2022). Consistent with the gradual phase-in of impacts, 31 respondents (60%) reported no adults had skipped meals in the last week. The remaining respondents provided explanations for persistent needs consistent with a wide range of potential economic factors including larger-than-average needs, inter-temporal substitution, inter-household risk-sharing, and non-consumption substitution. 9 respondents (17%) indicated the payment size was insufficient for everyone in their household to eat at least bread and tea at each meal, with several highlighting the difficulty of feeding larger families. 12 respondents (23%) indicated saving some food for future consumption, but typically only mentioned having one or two weeks’ worth on hand. 4 respondents (8%) reported sharing food with individuals outside their household, and 3 respondents (6%) reported bartering food for other expenses like medicine or rent.
@@ 1 Cost-efficiency
Table A13 reports raw costs and estimated cost-efficiency quantities. Panel A reports raw costs (a) disaggregated by category, (b) including/excluding onboarding, and (c) by treatment and control group status, which both received transfers but at different times. Panel B reports the CPB, including/excluding the transfer value, and including/excluding onboarding. The total CPB is $192.00$ USD consisting of four approximately $45$ USD transfers amounting to $180$ USD and the variable operations costs totaling $12$ USD. Excluding the costs of onboarding, the total CPB is $182.44$ USD. Here, the transfer amount is the same ($180$ USD) but the administrative costs are $2.44$ USD. Panel C reports the TCTR and CTR, first including onboarding, and then excluding onboarding costs. When including all costs, the TCTR is $1.067$, meaning that $6.7$ cents is required for each dollar transferred to a beneficiary (CTR). When excluding the costs of onboarding, the TCTR is $1.014$, meaning that $1.4$ cents is required for each dollar transferred (CTR).
@@ 2 Cost-effectiveness
We compute cost-effectiveness ratios (CERs) by dividing costs by treatment effects aggregated over participants and time (Riegg Cellini and Edwin Kee, 2015), which are interpreted as the number of dollars required to achieve a defined impact. All CER estimates appear in Table A14.
@@
Because digital aid continued to have an effect after the program ended (see Figure A2), CERs can reasonably include longer-run effects in which costs remain the same, but are now scaled by the cumulative reduction in days with skipped meals across all sixteen weeks. For the post-program period, the treatment group is compared to the control group’s baseline value (2.615) because in weeks 9–16 the control group is now receiving transfers and no longer constitutes a control. The treatment group’s 12- and 16-week values could be compared to those of the control group just before receiving treatment (2.546), but by that point the control group could have experienced other changes, including updated expectations about entering the program or any effect of receiving the monthly 350 AFA survey incentive, which almost all households received given the high response rates. Alternatively, they could be compared to the treatment group’s own baseline value (2.626) or an average of the treatment and control groups’ baseline values. In practice, all three are similar and do not qualitatively change the inferences. Based on these inputs and assumptions, the approximately $180$ USD in digital aid to 1208 beneficiaries resulted in a reduction of 15,844.13 days that households skipped meals (7,344.64 during the eight-week program period and 8,499.49 in the eight-week post-program period) out of a possible $135,296$ possible days across the sixteen weeks. This is a 11.71 percentage point reduction relative to not receiving assistance, making 14.64 USD the cost to reduce a day with skipped meals (\$231,936.08/15,844.13). The corresponding, cumulative calorie intake estimate is 209,950,567 fewer forgone calories for the treatment group (15,844.13 fewer days * 13,251 calories per household-day), making 1.10 USD the cost to increase calorie intake by 1,000 (\$231,936.08/(209,950,567/1000)).
@@ Appendix E Supplementary material figures and tables
(a) Skipped Meals
(b) Total Bread and Tea Meals
(c) Life Satisfaction
(d) Happy
Figure A2: Effects over Entire Duration of Program
@@
(1) (2) (3) (4) (5) Panel A. Types of Food Rice Beans Vegetables Chicken Dairy $\beta_{1}$: Treated $\times$ Round 2 0.367*** 0.240*** 0.014 0.001 0.032 (0.060) (0.049) (0.071) (0.011) (0.023) $\beta_{2}$: Round 2 –0.051 –0.124*** 0.165*** –0.008 –0.065*** (0.040) (0.030) (0.051) (0.006) (0.015) $\beta_{3}$: Treated 0.412*** 0.373*** –0.010 0.012 0.031 (0.049) (0.041) (0.056) (0.009) (0.021) Control Mean 0.698 0.518 1.394 0.021 0.074 Observations 4,763 4,763 4,763 4,763 4,763 R Squared 0.083 0.080 0.010 0.011 0.012 p-value $\beta_{1}$ + $\beta_{3}$ = 0 0 0 0.942 0.046 0 Panel B. Other Economic Outcomes Income Employed Fin. Decision-Maker Medicine Purchase $\beta_{1}$: Treated $\times$ Round 2 36.807 0.050** 0.013 0.008 (151.022) (0.023) (0.022) (0.019) $\beta_{2}$: Round 2 –165.446** 0.032* 0.030** –0.046*** (64.257) (0.017) (0.015) (0.012) $\beta_{3}$: Treated 110.705 –0.033* 0.010 0.031** (149.712) (0.018) (0.019) (0.015) Control Mean 876.683 0.192 0.666 0.051 Observations 4,763 4,741 4,757 3,582 R Squared 0.010 0.021 0.061 0.011 p-value $\beta_{1}$ + $\beta_{3}$ = 0 0.048 0.361 0.202 0.002
@@
Experimenter Demand Control Baseline Overall Control Treatment Mean Estimate Estimate Estimate Estimate Variable (1) (2) (3) (4) (5) Panel A. Food Security Days skipping meals (past week) 2.615 –0.76*** –0.027 –0.099 0.039 (0.051) (0.057) (0.079) (0.068) Children skipping meals (=1) 0.873 –0.117*** 0.016 0.027 0.005 (0.012) (0.015) (0.018) (0.023) Regularly eat twice a day 0.501 0.093*** –0.005 –0.006 –0.003 (0.015) (0.013) (0.018) (0.019) Total bread and tea meals (past week) 13.639 –1.608*** –0.181 –0.107 –0.268 (0.121) (0.146) (0.199) (0.19) Food Security - KLK Index 0 0.501*** 0.002 –0.002 0.01 (0.032) (0.036) (0.045) (0.047) Panel B. Informal Taxation Inf. tax gov. off. (others) 0.004 0.001 –0.001 –0.002 0 (0.001) (0.002) (0.002) (0.003) Inf. tax comm. leader (others) 0.001 0.002 –0.002 0.002 –0.006 (0.002) (0.002) (0.002) (0.005) Inf. tax gov. off. (you) 0 0.002** –0.001 0 –0.002 (0.001) (0.001) (0) (0.003) Inf. tax comm. leader (you) 0.003 0.002 –0.005* –0.005* –0.005 (0.003) (0.003) (0.003) (0.004) Informal tax. - KLK Index 0 0.074* –0.077 –0.027 –0.124 (0.042) (0.054) (0.036) (0.106) Panel C. Wellbeing Better economic situation 0.048 0.335*** 0.003 0.018 –0.009 (0.011) (0.017) (0.014) (0.025) Satisfied with fin. situation 0.133 0.263*** 0.003 0.013 –0.005 (0.012) (0.017) (0.018) (0.024) Happy 0.009 0.28*** –0.007 –0.007 –0.004 (0.014) (0.016) (0.018) (0.021) Life satisfaction (std) 0.011 1.682*** –0.007 0.073 –0.075 (0.058) (0.079) (0.085) (0.103) Economic/Wellbeing - KLK Index 0 1.498*** 0.001 0.054 –0.04 (0.042) (0.063) (0.055) (0.081)
@@
$Notes$: Controls for stratification fixed effects, survey round fixed effects (baseline estimates), and baseline values of the dependent variable, if available, are included. Standard errors clustered at the individual level are in parentheses. Control mean is the mean in the baseline if available, or across follow up rounds otherwise, for the control group. The KLK Index is created following Katz, Kling, & Liebman (2007), and is the equally-weighted sum of the standardised component variables. The baseline effect is the (pooled) ITT effect of the main treatment (receiving the aid payments). The overall effect is the coefficient on the prime treatment. The control effect is the coefficient on the prime treatment in a regression where the prime treatment and the main treatment are interacted, while the effect on the treated is the sum of the prime treatment and the interaction term between the two treatments from the same regression. Better economic situation is an index that equals 1 if the respondent answered that her economic situation compared to 30 days ago is slightly or much better, and 0 otherwise. Satisfied with financial situation is a dummy that equals 1 if the respondent answered that she agrees a lot or somewhat with the statement that she is highly satisfied with her current financial condition, and 0 otherwise. Happy is a dummy that equals 1 if respondent said that she was very happy or quite happy, and 0 otherwise. Life satisfaction is the score from 1 (dissatisfied) to 10 (satisfied) in terms of how satisfied the respondent is with her life as a whole these days (standardised). Total household income excludes the aid payments.
@@
Control Control Treatment Standard Naive Adjusted Mean SD Effect Error p-value p-value N (1) (2) (3) (4) (5) (6) (7) Panel A. Primary Outcomes Days skipping meals (past week) 2.579 1.813 –0.785 0.053 0 0.0002 4412 Children skipping meals (=1) 0.87 0.336 –0.113 0.012 0 0.0002 4412 Regularly eat twice a day 0.493 0.5 0.1 0.015 0 0.0002 4412 Total bread and tea meals (past week) 13.644 3.921 –1.628 0.125 0 0.0002 4412 Food Security - KLK Index –0.016 1.01 0.51 0.033 0 4412 Food Security - Anderson Index 0.002 1.058 0.493 0.034 0 4412 Inf. tax gov. off. (others) 0.002 0.043 0.001 0.001 0.563 0.6877 4412 Inf. tax comm. leader (others) 0.002 0.043 0.003 0.002 0.118 0.1216 4412 Inf. tax gov. off. (you) 0.001 0.03 0.002 0.001 0.163 0.151 4412 Inf. tax comm. leader (you) 0.005 0.071 0.001 0.002 0.612 0.6877 4412 Informal tax. - KLK Index –0.007 0.923 0.075 0.041 0.068 4412 Informal tax. - Anderson Index –0.016 0.763 0.048 0.031 0.116 4412 Better economic situation 0.047 0.212 0.34 0.011 0 0.0002 4412 Satisfied with fin. situation 0.133 0.339 0.264 0.013 0 0.0002 4412 Happy 0.235 0.424 0.275 0.014 0 0.0002 4412 Life satisfaction 2.969 1.888 1.999 0.07 0 0.0002 4412 Economic/Wellbeing - KLK Index 0.001 1 1.509 0.044 0 4412 Economic/Wellbeing - Anderson Index 0.096 1.167 1.326 0.045 0 4412 Panel B. Secondary Outcomes Days eating rice (past week) 0.672 1.05 0.615 0.036 0 0.001 4412 Days eating beans (past week) 0.517 0.841 0.491 0.03 0 0.001 4412 Days eating vegetables (past week) 1.377 1.331 0 0.042 0.995 0.42 4412 Days eating chicken (past week) 0.018 0.138 0.014 0.006 0.021 0.022 4412 Days eating dairy (past week) 0.073 0.387 0.05 0.014 0 0.001 4412 Able to buy medicine 0.078 0.269 0.031 0.01 0.003 0.005 3344 Involved in fin. decisions 0.672 0.47 0.019 0.016 0.23 0.152 4412 Total household income (past month) 1093.634 1792.27 156.559 94.73 0.099 0.071 4412 Household’s head employed (past month) 0.284 0.451 –0.008 0.014 0.579 0.278 4412
@@
Table A6: Are Digital Payments Diverted? Alternative Indices
KLK Index KLK Index Anderson Index KLK Dropping Others You All Outlier (1) (2) (3) (4) Panel A. Baseline Treated 0.034 0.043 0.048 0.059 (0.026) (0.031) (0.036) (0.039) Observations 4,611 4,647 4,648 4,647 Control Mean –0.003 0.015 0.011 0.012 Panel B. Long-Run $\beta_{1}$: Treated $\times$ Round 2 0.073 0.025 0.052 0.063 (0.052) (0.061) (0.071) (0.079) $\beta_{2}$: Round 2 –0.048* –0.097*** –0.135*** –0.137*** (0.029) (0.036) (0.046) (0.045) $\beta_{3}$: Treated –0.074 0.006 –0.029 –0.035 (0.084) (0.109) (0.125) (0.136) Observations 4,611 4,647 4,648 4,647 ($\beta_{1}$ + $\beta_{2}$ + $\beta_{3}$) / $\beta_{3}$ 0.670 –11.440 3.920 3.120 p-value $\beta_{1}$ + $\beta_{3}$ = 0 0.964 0.556 0.701 0.663
@@
Table A8: Actual Values vs. Experts’ Predictions
Variable Actual Values Predicted Values p-value Predicted = Actual (1) (2) (3) How Many Bread and Tea Meals in Past Week? 11.96 10.32 $<$ 0.001 (0.45) Percent Reporting Diversion Attempts 1.99 39.85 $<$ 0.001 (2.98) Percent Able to Use Digital Payments 99.75 43.82 $<$ 0.001 (2.75) Delivery Costs 6.7 10.65 $<$ 0.001 (0.74)

### v4

@@ 5 Interpreting magnitudes of needs results
The Afghanistan Cash and Voucher Working Group (CVWG) estimates a Basic Food Basket for a family of seven costs approximately 96 USD per month (Bete, 2022), roughly equivalent to two biweekly direct aid transfers of 4000 AFA and a monthly survey incentive of 350 AFA. This basket was composed of 89 kg wheat flour, 21 kg domestic rice, 7 liters vegetable oil, 9 kg puSlses, and 1 kg salt at prevailing exchange rates in August 2022. For reference, the Minimum Expenditure Basket including food, healthcare, shelter, and other components totaled almost twice as much at 181.36 USD. In mid-December 2022, we contacted 25 merchants serving beneficiaries to better understand spending patterns and realized prices. We confirmed that the most popular purchases were wheat flour, cooking oil and sugar, and using merchant-specific estimates of prices and volumes, verified that the cost of a typical basket of goods matched the total value of the aid payments and survey incentive.
It is natural to ask whether the estimated impacts on basic needs are in line with what we should expect given the size of payments – and more specifically, why hunger appears to persist throughout the length of the intervention. As noted above, we observe large reductions in each of the four primary food security outcomes, and these measures kept improving throughout the two months of payments (Table 2, Panel A). While not eliminated, the number of days in which a person skipped meals in the past week declined after each payment (Figure 2, Panel A). To better understand the explanatory factors for this persistence, we completed a brief survey on January 2 & 3, 2023 with 52 randomly selected treatment group participants who reported any skipped meals in the second month of surveys (ie. December 2022). Consistent with the gradual phase-in of impacts, 31 respondents (60%) reported no adults had skipped meals in the last week. The remaining respondents provided explanations for persistent needs consistent with a wide range of potential economic factors including larger-than-average needs, inter-temporal substitution, inter-household risk-sharing, and non-consumption substitution. 9 respondents (17%) indicated the payment size was insufficient for everyone in their household to eat at least bread and tea at each meal, with several highlighting the difficulty of feeding larger families. 12 respondents (23%) indicated saving some food for future consumption, but typically only mentioned having one or two weeks’ worth on hand. 4 respondents (8%) reported sharing food with individuals outside their household, and 3 respondents (6%) reported bartering food for other expenses like medicine or rent.
@@ 1 Cost-efficiency
A13 reports raw costs and estimated cost-efficiency quantities. Panel A reports raw costs (a) disaggregated by category, (b) including/excluding onboarding, and (c) by treatment and control group status, which both received transfers but at different times. Panel B reports the CPB, including/excluding the transfer value, and including/excluding onboarding. The total CPB is $192.00$ USD consisting of four approximately $45$ USD transfers amounting to $180$ USD and the variable operations costs totaling $12$ USD. Excluding the costs of onboarding, the total CPB is $182.44$ USD. Here, the transfer amount is the same ($180$ USD) but the administrative costs are $2.44$ USD. Panel C reports the TCTR and CTR, first including onboarding, and then excluding onboarding costs. When including all costs, the TCTR is $1.067$, meaning that $6.7$ cents is required for each dollar transferred to a beneficiary (CTR). When excluding the costs of onboarding, the TCTR is $1.014$, meaning that $1.4$ cents is required for each dollar transferred (CTR).
@@ 2 Cost-effectiveness
We compute cost-effectiveness ratios (CERs) by dividing costs by treatment effects aggregated over participants and time (Riegg Cellini and Edwin Kee, 2015), which are interpreted as the number of dollars required to achieve a defined impact. All CER estimates appear in A14.
@@
Because digital aid continued to have an effect after the program ended (see Figure 2), CERs can reasonably include longer-run effects in which costs remain the same, but are now scaled by the cumulative reduction in days with skipped meals across all sixteen weeks. For the post-program period, the treatment group is compared to the control group’s baseline value (2.615) because in weeks 9–16 the control group is now receiving transfers and no longer constitutes a control. The treatment group’s 12- and 16-week values could be compared to those of the control group just before receiving treatment (2.546), but by that point the control group could have experienced other changes, including updated expectations about entering the program or any effect of receiving the monthly 350 AFA survey incentive, which almost all households received given the high response rates. Alternatively, they could be compared to the treatment group’s own baseline value (2.626) or an average of the treatment and control groups’ baseline values. In practice, all three are similar and do not qualitatively change the inferences. Based on these inputs and assumptions, the approximately $180$ USD in digital aid to 1208 beneficiaries resulted in a reduction of 15,844.13 days that households skipped meals (7,344.64 during the eight-week program period and 8,499.49 in the eight-week post-program period) out of a possible $135,296$ possible days across the sixteen weeks. This is a 11.71 percentage point reduction relative to not receiving assistance, making 14.64 USD the cost to reduce a day with skipped meals (\$231,936.08/15,844.13). The corresponding, cumulative calorie intake estimate is 209,950,567 fewer forgone calories for the treatment group (15,844.13 fewer days * 13,251 calories per household-day), making 1.10 USD the cost to increase calorie intake by 1,000 (\$231,936.08/(209,950,567/1000)).
@@ Appendix E Supplementary material figures and tables
Table A2: Baseline Balance Check
Whole Sample Treatment Control p-value Mean SD Mean Mean Difference Variable (1) (2) (3) (4) (5) Panel A. Outcome Vars 1. Total bread-tea meals 13.76 2.56 13.74 13.77 0.79 2. Skipped meals 2.62 1.52 2.63 2.62 0.86 3. Inf tax gov (others) 0.003 0.058 0.002 0.004 0.474 4. Inf tax leaders (others) 0 0.02 0 0.001 0.317 5. Inf tax gov (you) 0 0 0 0 1 6. Inf tax leaders (you) 0.002 0.041 0 0.003 0.045 7. Life satisfaction 3.53 1.17 3.51 3.54 0.59 8. Happy 0.01 0.1 0.01 0.01 0.99 Panel B. Heterogeneity Vars 1. Married 0.34 0.47 0.34 0.34 0.88 2. Pashtun 0.1 0.3 0.11 0.09 0.15 3. Some education 0.37 0.48 0.37 0.36 0.37 4. Balkh 0.33 0.47 0.33 0.33 0.77 5. Herat 0.33 0.47 0.33 0.33 0.85 6. Kabul 0.34 0.47 0.34 0.34 0.95 7. Above median age 0.56 0.5 0.56 0.56 0.81 8. Fin. decision maker 0.66 0.47 0.66 0.66 0.99 9. Above median HH size 0.64 0.48 0.64 0.65 0.6 10. Able to leave house 0.65 0.48 0.64 0.66 0.56 Number of individuals 2409 1208 1201
$Notes$: The table shows, for different pre-specified variables at baseline, the overall mean and standard deviation (columns 1 and 2), the mean in the treatment group (column 3) and the mean in the control group (column 4). Column 5 shows the p-value of the difference between the means in the treatment and control group, adjusting for robust standard errors.
@@
(1) (2) (3) (4) (5) Panel A. Types of Food Rice Beans Vegetables Chicken Dairy $\beta_{1}$: Treated $\times$ Round 2 0.366*** 0.240*** 0.014 0.001 0.032 (0.061) (0.050) (0.071) (0.011) (0.023) $\beta_{2}$: Round 2 –0.051 –0.124*** 0.164*** –0.008 –0.065*** (0.040) (0.030) (0.051) (0.006) (0.015) $\beta_{3}$: Treated 0.413*** 0.373*** –0.011 0.012 0.031 (0.049) (0.041) (0.056) (0.009) (0.021) Control Mean 0.698 0.518 1.394 0.021 0.074 Observations 4,763 4,763 4,763 4,763 4,763 R Squared 0.088 0.084 0.013 0.012 0.014 Panel B. Other Economic Outcomes Income Employed Fin. Decision-Maker Medicine Purchase $\beta_{1}$: Treated $\times$ Round 2 36.972 0.050** 0.014 0.008 (151.108) (0.023) (0.022) (0.019) $\beta_{2}$: Round 2 –165.374** 0.032* 0.030** –0.046*** (64.319) (0.017) (0.015) (0.012) $\beta_{3}$: Treated 109.985 –0.033* 0.010 0.031** (149.582) (0.018) (0.019) (0.015) Control Mean 876.683 0.192 0.666 0.051 Observations 4,763 4,741 4,757 3,582 R Squared 0.012 0.024 0.064 0.015
@@
Experimenter Demand Control Baseline Overall FWER Control Treatment Mean Estimate Estimate $p$-value Estimate Estimate Variable (1) (2) (3) (4) (5) (6) Panel A. Food Security Days skipping meals (past week) 2.615 –0.76*** –0.031 0.7225 –0.082 0.041 (0.051) (0.062) (0.087) (0.077) Children skipping meals (=1) 0.873 –0.117*** 0.016 0.3913 0.028 0.009 (0.012) (0.016) (0.021) (0.024) Regularly eat twice a day 0.501 0.093*** –0.001 0.925 –0.004 –0.001 (0.015) (0.014) (0.02) (0.02) Total bread and tea meals (past week) 13.639 –1.608*** –0.186 0.3897 –0.059 –0.267 (0.121) (0.161) (0.212) (0.217) Food Security - KLK Index 0 0.501*** 0.005 –0.011 0.007 (0.032) (0.039) (0.05) (0.051) Panel B. Informal Taxation Inf. tax gov. off. (others) 0.004 0.001 –0.002 0.4468 –0.002 –0.001 (0.001) (0.002) (0.002) (0.003) Inf. tax comm. leader (others) 0.001 0.002 –0.002 0.4513 0 –0.005 (0.002) (0.003) (0.002) (0.005) Inf. tax gov. off. (you) 0 0.002** –0.001 0.4513 –0.001 –0.001 (0.001) (0.001) (0.001) (0.003) Inf. tax comm. leader (you) 0.003 0.002 –0.006** 0.2026 –0.006 –0.006 (0.003) (0.003) (0.004) (0.004) Informal tax. - KLK Index 0 0.074* –0.1* –0.075 –0.127 (0.042) (0.057) (0.05) (0.093) Panel C. Wellbeing Better economic situation 0.048 0.335*** 0.006 0.9754 0.016 –0.012 (0.011) (0.019) (0.017) (0.026) Satisfied with fin. situation 0.133 0.263*** –0.003 0.9754 0.005 –0.017 (0.012) (0.018) (0.02) (0.026) Happy 0.009 0.28*** –0.005 0.9754 –0.014 –0.003 (0.014) (0.017) (0.02) (0.022) Life satisfaction (std) 0.011 1.682*** 0.002 0.9754 0.043 –0.081 (0.058) (0.089) (0.098) (0.111) Economic/Wellbeing - KLK Index 0 1.498*** 0.004 0.027 –0.059 (0.042) (0.071) (0.064) (0.086)
@@
$Notes$: Control for stratification fixed effects, survey round fixed effects (baseline estimates), and baseline value of dependent variable, if available. Standard errors clustered at individual level in parenthesis. Control mean is the mean in the baseline if available, or across follow up rounds otherwise, for the control group. The KLK Index is created following Katz, Kling, & Liebman (2007), and is the equally-weighted sum of the standardised component variables. The baseline effect is the (pooled) ITT effect of the main treatment (receiving the aid payments). Primary outcomes show FWER-adjusted p-values within each family outcome (following Romano & Wolf, 2005, using 5000 repetitions). The overall effect is the coefficient on the prime treatment. The control effect is the coefficient on the prime treatment in a regression where the prime treatment and the main treatment are interacted, while the effect on the treated is the sum of the prime treatment and the interaction term between the two treatments from the same regression. Better economic situation is an index that equals 1 if the respondent answered that her economic situation compared to 30 days ago is slightly or much better, and 0 otherwise. Satisfied with financial situation is a dummy that equals 1 if the respondent answered that she agrees a lot or somewhat with the statement that she is highly satisfied with her current financial condition, and 0 otherwise. Happy is a dummy that equals 1 if respondent said that she was very happy or quite happy, and 0 otherwise. Life satisfaction is the score from 1 (dissatisfied) to 10 (satisfied) in terms of how satisfied the respondent is with her life as a whole these days (standardised). Total household income excludes the aid payments.
Levels of significance: *$p<0.1$ , **$p<0.05$ , ***$p<0.01$.
@@
Control Control Treatment Standard Naive Adjusted Mean SD Effect Error p-value p-value N (1) (2) (3) (4) (5) (6) (7) Panel A. Primary Outcomes Days skipping meals (past week) 2.579 1.813 –0.785 0.053 0 0.0002 4412 Children skipping meals (=1) 0.87 0.336 –0.113 0.012 0 0.0002 4412 Regularly eat twice a day 0.493 0.5 0.1 0.015 0 0.0002 4412 Total bread and tea meals (past week) 13.644 3.921 –1.628 0.124 0 0.0002 4412 Food Security - KLK Index –0.016 1.01 0.511 0.033 0 4412 Food Security - Anderson Index 0.002 1.058 0.494 0.034 0 4412 Inf. tax gov. off. (others) 0.002 0.043 0.001 0.001 0.563 0.6818 4412 Inf. tax comm. leader (others) 0.002 0.043 0.003 0.002 0.116 0.108 4412 Inf. tax gov. off. (you) 0.001 0.03 0.002 0.001 0.163 0.1524 4412 Inf. tax comm. leader (you) 0.005 0.071 0.001 0.002 0.609 0.6818 4412 Informal tax. - KLK Index –0.007 0.923 0.075 0.041 0.067 4412 Informal tax. - Anderson Index –0.016 0.763 0.048 0.031 0.115 4412 Better economic situation 0.047 0.212 0.34 0.011 0 0.0002 4412 Satisfied with fin. situation 0.133 0.339 0.264 0.013 0 0.0002 4412 Happy 0.235 0.424 0.275 0.014 0 0.0002 4412 Life satisfaction 2.969 1.888 2 0.07 0 0.0002 4412 Economic/Wellbeing - KLK Index 0.001 1 1.51 0.043 0 4412 Economic/Wellbeing - Anderson Index 0.096 1.167 1.326 0.045 0 4412 Panel B. Secondary Outcomes Days eating rice (past week) 0.672 1.05 0.616 0.036 0 0.001 4412 Days eating beans (past week) 0.517 0.841 0.491 0.03 0 0.001 4412 Days eating vegetables (past week) 1.377 1.331 0 0.042 0.991 0.415 4412 Days eating chicken (past week) 0.018 0.138 0.014 0.006 0.021 0.022 4412 Days eating dairy (past week) 0.073 0.387 0.05 0.014 0 0.001 4412 Able to buy medicine 0.078 0.269 0.031 0.01 0.002 0.004 3344 Involved in fin. decisions 0.672 0.47 0.019 0.016 0.228 0.15 4412 Total household income (past month) 1093.634 1792.27 156.196 94.148 0.097 0.07 4412 Household’s head employed (past month) 0.284 0.451 –0.008 0.014 0.573 0.274 4412
@@
Table A7: Are Digital Payments Diverted? Indices
KLK Index KLK Index Anderson Index Dropping Others You All Outlier (1) (2) (3) (4) Panel A. Baseline Treated 0.034 0.043 0.049 0.059 (0.026) (0.031) (0.036) (0.039) Observations 4,611 4,647 4,648 4,647 Control Mean –0.003 0.015 0.011 0.012 Panel B. Long-Run $\beta_{1}$: Treated $\times$ Round 2 0.072 0.024 0.051 0.062 (0.053) (0.061) (0.071) (0.079) $\beta_{2}$: Round 2 –0.048 –0.096*** –0.135*** –0.137*** (0.029) (0.037) (0.046) (0.045) $\beta_{3}$: Treated –0.073 0.007 –0.027 –0.034 (0.085) (0.109) (0.125) (0.137) Observations 4,611 4,647 4,648 4,647 ($\beta_{1}$ + $\beta_{2}$ + $\beta_{3}$) / $\beta_{3}$ 0.670 –9.410 4.090 3.230